### v1

@@ 7.3 Japanese vowels : non-stationary classification
Another speech recognition task uses a dataset [^9] Jaeger et al., 2007a give two links where the dataset can be found. Those links no longer work; the dataset can be found at https://github.com/MaterialMan/CHARC/tree/master/Support%20files/other/Datasets/JapaneseVowels. of vowels spoken by nine Japanese men (Jaeger et al., 2007a, ). Unlike the spoken digits task, the task is not to categorise by what has been spoken, but rather to categorise by speaker. Hence it remains a multiple-output classification task. In the original paper, 4 experimental setups were tested, the most successful of which we summarise here.
The experiment used 9 output nodes, one for each speaker, and had a training length of 270 samples.
The inputs were processed by first being partitioned into $D=3$ subsequences of equal length; these subsequences were then joined into single input vectors of size $D$. The outputs were then trained using linear regression. Because overfitting was a concern, the reservoir size was kept to $N=4$.
@@ 7.4 Santa Fe sleep apnea dataset: stationary classification
Figure 15: Portion of the sleep apnea dataset, as used in Wringe et al., (2024); each dataset portion is separately normalised to the range $[-0.5,0.5]$ using equation 20.
The Santa Fe 1992 Time Series Prediction competition sleep apnea dataset (Ichimaru and Moody,, 1999), available at PhysioNet (Goldberger et al.,, 2000) comprises various physiological readings of a patient with suspected sleep apnea. The patient experienced both normal sleep and apnea periods, which can be seen in the data (figure 14).
The data was gathered over 4 hours and 43 minutes with readings taken every 0.5 seconds, then digitised at 250 Hz. It is part of the Santa Fe Time Series Prediction Competition (Schwenker and Labib,, 2009), in two files, b1.txt and b2.txt, each with 17000 datapoints, with an accompanying file describing the content. The dataset consists of readings of the patient’s (i) heartrate, (ii) respiration rate, and (iii) blood oxygen saturation, although there is conflicting information as to whether the respiration data is derived from the patient’s nasal airflow (Schwenker and Labib,, 2009) or chest volume (Ichimaru and Moody,, 1999).
@@
We designate this as a stationary temporal classification task, as it could potentially be used to classify normal v apnea periods. However, in the literature, the dataset tends to be used for a time series prediction task, either predicting each datastream individually, or using a combination to improve prediction. For example, Shougat et al., (2021), who treat each of the datasets as entirely distinct benchmark tasks, while Wringe et al., (2024) study how the datasets interact with each other.
Shougat et al., (2021) state that the input is normalised before being input to the reservoir, but it is unclear how this normalisation is performed. Wringe et al., (2024) normalise the inputs to fall between $[-0.5,0.5]$ using:
@@
The data lengths used by Shougat et al., (2021) are not stated. Wringe et al., (2024) start their input at point 22000 (fig. 15), and use 1000 points as washout, 3000 points as training, and 1000 points as testing. These values are chosen to ensure both the training and testing phases include both normal sleep and apnea periods.
7.5 McMaster IPIX radar data : unknown dynamics
The McMaster Ipix Radar is an X-band radar, originally created to detect icebergs, but capable of collecting data on all sea clutter. The website (Bakker and Currie,, 2001) indicates that the movement of sea clutter is influenced by several dynamical systems, and is itself a nonlinear dynamical system. As such, it makes an ideal prediction benchmark for Reservoir Computing.
The radar’s website makes two databases available for general use (Bakker and Currie,, 2001), one created from data recorded in Dartmouth, Nova Scotia, in 1993, and the other from data collected in Grimsby, Ontario, in 1998. The latter database is marked as incomplete by the radar’s website, and is encouraged to be used only comparatively with the Dartmouth data.
This data has been used as a prediction task in Reservoir Computing (Xue et al.,, 2007). It is unclear exactly which of the datasets are used, or how these are provided as input for the Reservoir Computer. 2000 datapoints are used, of which 200 are used as washout data, 800 for training, and 1000 for testing, but no more information is given.
Subsequent uses of this dataset in Reservoir Computing (Čerňanský and Tiňo,, 2008; Rodan and Tiňo,, 2010; Rodan and Tino,, 2011; Duport et al., 2012a, ; Dejonkheere et al.,, 2014) appear to be consistent with the first use, citing (Xue et al.,, 2007) and using the same dataset sizes.
@@ 8.1.1 Linear Memory Capacity
Certain tasks are interesting because they directly reveal properties of a given Reservoir Computer. One of these is the Linear Memory Capacity task, which quantifies the fading memory of the reservoir. This was introduced by Jaeger, (2002), and has been investigated in the context of reservoir computing and used as a benchmark task (Verstraeten et al.,, 2007; Rodan and Tiňo,, 2010; Duport et al., 2012a, ; Goudarzi et al.,, 2015; Soriano et al.,, 2015).
The Linear Memory Capacity of a given Reservoir Computer that takes scalar input is defined as follows. Consider an input stream $u(t)\in U[-1,1]$. Train the reservoir to reproduce this input, delayed by a number of timesteps $k$: the target output is ${\hat{v}}=u(t-k)$, the observed output is $v_{k}(t)$, the linear memory capacity for this delay is defined as the covariance squared of the delayed input (target output) and the observed output, normalised by the variances of the input and the observed output:
@@
The total Linear Memory Capacity is the sum over all delays:
@@
For higher $k$, each $MC_{k}$ tends to decrease (and the corresponding validation NRMSE tends to increase), which means that inputs after longer delays are remembered less well, and are dominated by noise. In the formal definition of $MC$ (eqn. 22), the sum over delays goes to infinity; however the small values at high $k$ are essentially noise, and should be neglected, so in practice a cutoff is used. Dambre et al., (2012, SupMat3.2) define a threshold based on the size of the reported capacity. Dale, 2018b uses a cutoff of $k_{max}=2N$ (number of nodes).
@@ 8.1.2 Nonlinear Memory Capacity
While the Linear Memory Capacity measure gives a useful quantification of a reservoir’s memory capacity, it cannot on its own model the full computing power of dynamical systems. (Dambre et al.,, 2012) generalise the definition of linear memory capacity in a way that allows them to define various non-linear capacities, too: they call this information processing capacity (IPC). These measure a Reservoir Computer’s ability to compute a nonlinear function of past inputs, for example, a cubic function of delayed inputs, such as $u^{3}(t-1)$ or $u(t-1)u^{2}(t-2)$. The cases where the polynomials involve different time delays are called cross memory capacities (Duport et al., 2012b, ).
Duport et al., 2012b consider quadratic polynomials. Dambre et al., (2012) consider complete sets of orthogonal polynomials, to cover all possible non-linear capacities and cross memory capacities with no double counting. They use normalised Legendre polynomials for inputs $u(t)\in[-1,1]$. They note that different sets of orthogonal polynomials should be used for different input distributions, for example, Hermite polynomials for Gaussian-distributed inputs. Orthogonal sets of trigonometric functions can also be used as the basis.
As for linear memory capacity, the reservoir is trained to output the relevant (here, polynomial) function of its delayed input. For a given degree $d$ of polynomial (linear, quadratic, cubic, etc), the contributions for all delays $k$, including all combinations of delays in the cross memory capacities, are summed. Then the total non-linear memory capacity, or IPC, is the sum over all polynomial degrees. Dambre et al., (2012) prove that the total IPC (from the contributions of all the linear and non-linear polynomials) is $MC=N$.
@@ 8.2 Rank-based measures
Legenstein and Maass, (2007); Büsing et al., (2010) introduce two particular rank-based measures: kernel quality (kernel rank, KR), and generalisation rank (GR). These have been adapted for use with ESNs in two different ways in the literature, discussed below. Calculating the rank is discussed in section 8.2.3.
@@ 8.2.1 Kernel Rank
Vidamour et al., (2022) use a direct translation of the original definitions, and implement this measure as follows. Consider $S$ maximally distinct input streams, with values drawn from $U[-1,1]$, each of length $T$, and measure the reservoir state at the end of each input stream. KR is the rank of the resulting $N\times S$ matrix $[\mathbf{x}_{1}\mathbf{x}_{2}\ldots\mathbf{x}_{S}]$, see algorithm 1.
Algorithm 1 KR, Vidamour et al., (2022)
@@
Dale et al., (2019) use a different adaptation, which is computationally less intensive (requiring $S$ inputs, rather than $S\times T$), and adapted to time series tasks (where the reservoir state at each timepoint is used), but is further from the original definitions. In this approach, there is a single input stream, of length $S$, and the reservoir state is measured at each timepoint. The stream’s values are again drawn from $U[-1,1]$, thereby making them maximally distinct. See algorithm 2.
Algorithm 2 KR and GR, Dale et al., (2019)
@@ 8.2.2 Generalisation Rank
Vidamour et al., (2022) again use a direct translation of the original definitions. Consider $S$ input streams, each of length $T$, each with values drawn from $U[-1,1]$, except that the last few values in each stream are set to be the same for all streams. Measure the reservoir state at the end of each input stream. GR is the rank of the resulting matrix, see algorithm 3.
Algorithm 3 GR, Vidamour et al., (2022)
@@
Dale et al., (2019) again use a different adaptation of the original definitions. The algorithm for GR is the same as for KR, except that the streams’ values are drawn from a reduced range $U[-0.1,0.1]$, thereby making them similar. See algorithm 2.
@@ 8.2.3 Calculating the rank
An alternative to this integer-valued rank is the real-valued effective rank (Roy and Vetterli,, 2007; Love et al.,, 2021). Normalise the singular values $\sigma_{i}$ to sum to one: $p_{i}=\sigma_{i}/\sum_{i}\sigma_{i}$. The effective rank is defined as $\exp(-\sum_{i}p_{i}\ln p_{i})$. If $R$ of the singular values are the same, and the rest are zero, then $p_{i}=1/R$ or 0, giving an effective rank of $R$, which is the same as the standard rank value. For other cases, effective rank gives continuous values, has no arbitrary cutoff, and weights the singular values according to their size, potentially giving a more meaningful result.
The number of measured states $S$ should not be less than $N$, as the rank of the resulting matrix is $\leq\min(S,N)$. Dale et al., (2019, sec.D.a) note that the measured rank increases with $S$ until it eventually converges. Preliminary investigation should be performed to establish a suitable value for $S$.
@@ 8.3 Benchmarks for ‘free’
Certain tasks, such as MC and GR, can be calculated from the same data output from the ESN (Love et al.,, 2021). Both tasks use the same input; the observations of the state can be used to calculate the KR, and an output layer can be trained on the same observations to find the memory capacity. This is useful when using slow physical reservoirs, or computationally intensive search methods such as CHARC.
@@ 8.4 CHARC
Memory capacity, Kernel Rank, and Generalisation Rank are all relevant properties of Reservoir Computers, but the aim should not be to maximise them all. There is no agreed upon principle that allows us to correlate these values of these metric. Indeed, different tasks are best accomplished by reservoirs with different properties (Dale et al.,, 2019). Given this, they construct a framework based on Linear Memory Capacity, Kernel Rank, and Generalisation Rank. The Reservoir can then be evaluated not through its performance at the tasks used to measure these properties, but instead by the breadth of the behaviour space available to the substrate or model it is built upon.
@@
Dale et al., (2019) map four of the benchmark tasks described here to their positions in the CHARC behaviour space. One piece of interesting further work might be to map more of them, in order to enable studies to select tasks that not only take a range of different forms, but also correspond to different areas of Reservoir behaviour spaces.
@@ 9 Best practices for benchmarking
Having looked at many of the benchmarks used in the field of Reservoir Computing individually, we now draw out some best practices for using them.
@@
This section primarily concerns running and reporting benchmarks in reservoir computing. For a more general introduction and tutorial to performing RC, see Cucchi et al., (2022).
@@ 9.1.1 Using existing datasets
Some equation-based benchmarks, such as the NARMA (section 4.1 and Mackey–Glass (section 5.1 systems, rely on specific parameter values. Different values may lead to different behaviour, as with the Mackey–Glass system, where a $\tau$ of 16.8 or more will lead to chaotic behaviour. Sometimes several different sets of values are used in the literature.
@@ 9.1.2 Introducing new datasets
As noted by Gilpin, (2021), if a benchmark is based on a differential equation and numerically integrated, then the values used in the integration, such as timestep and gridsize, or the details of a more sophisticated algorithm, can change the detailed generated dataset.
@@
If using a new experimental or generated data set, it needs to be published with the results, and accompanied by metadata. Gebru et al., (2021) provide comprehensive guidelines for what metadata should be provided, covering issues including who created the dataset and who funded them, how and when the data was collected or generated, ethical concerns, the structure of the data, sampling, noise and errors, labelling, data cleaning and pre-processing, previous uses, distribution, and maintenance. Gilpin, (2021)’s catalogue of 131 chaotic dynamics benchmark systems follows the approach given there.
@@ 9.2.1 Reservoir specifications
The first aspect of experimental setup that needs to be reported are the parameter values of the Reservoir Computer used. One example of this being done well can be found by Jaeger, (2010), where the setup for the Mackey–Glass benchmark is described as having 400 nodes, bias input, inserted noise, and output feedback. When reporting an experiment involving an ESN, some parameters that should be reported are:
@@
A guide to setting these parameters can be found in Lukoševičius, (2012).
@@ 9.2.2 Washout, training, testing data
It is important to note the length of the washout set, particularly for common dataset-based benchmarks, where different washout lengths mean that the Reservoir Computer is being tested and trained on different subsets of the dataset. The training length of the reservoir can have an effect on its performance: up to a certain training length, increasing the training will lead to more consistent results across experiments, leading to a smaller standard deviation of different performances. After this saturation is reached, increasing the training length has little effect, and may result in overfitting. This saturation can be seen both in performance-based tasks such as NARMA, and value-based tasks such as Kernel Rank (Dale, 2018b, ).
@@
Similarly, different testing lengths may affect overall performance. For example, in a free-running prediction task (section 3.1.2), the experiment that uses a longer testing set may yield worse performance, due to the accumulation of errors leading to growing errors (Jaeger et al., 2007a, ).
9.2.3 Multiple runs

### v2

@@ 7.3 Japanese vowels : non-stationary classification
Another speech recognition task uses a dataset999[12] gives links to the dataset; those links no longer work. The dataset can currently be found at https://github.com/MaterialMan/CHARC/tree/master/Support%20files/other/Datasets/JapaneseVowels. of vowels spoken by nine Japanese men [12]. Unlike the spoken digits task, the task is not to categorise by what has been spoken, but rather to categorise by speaker. Hence it remains a multiple-output classification task.
In the original paper, four experimental setups were tested, the most successful of which we summarise here. The experiment used 9 output nodes, one for each speaker, and had a training length of 270 samples. The inputs were processed by first being partitioned into $D=3$ subsequences of equal length; these subsequences were then joined into single input vectors of size $D$. The outputs were then trained using linear regression. Because overfitting was a concern, the reservoir size was kept to $N=4$.
@@ 7.4 Santa Fe sleep apnea dataset: stationary classification
Figure 15: Portion of the sleep apnea dataset, as used in [120]; each dataset portion is separately normalised to the range $[-0.5,0.5]$ using equation 20.
The Santa Fe 1992 Time Series Prediction competition sleep apnea dataset [121], available at PhysioNet [122] comprises various physiological readings of a patient with suspected sleep apnea. The patient experienced both normal sleep and apnea periods, which can be seen in the data (figure 14).
The data was gathered over 4 hours and 43 minutes with readings taken every 0.5 seconds, then digitised at 250 Hz. It is part of the Santa Fe Time Series Prediction Competition [92], in two files, b1.txt and b2.txt, each with 17000 datapoints, with an accompanying file describing the content. The dataset consists of readings of the patient’s (i) heart rate, (ii) respiration rate, and (iii) blood oxygen saturation, although there is conflicting information as to whether the respiration data is derived from the patient’s nasal airflow [92] or chest volume [121].
@@
We designate this as a stationary temporal classification task, as it could potentially be used to classify normal v apnea periods. However, in the literature, the dataset tends to be used for a time series prediction task, either predicting each datastream individually, or using a combination to improve prediction. For example, [26] treats each of the datasets as entirely distinct benchmark tasks, while [120] studies how the datasets interact with each other.
Shougat et al. [26] state that the input is normalised before being input to the reservoir, but it is unclear how this normalisation is performed. Wringe [120] normalises the inputs to fall between $[-0.5,0.5]$ using:
@@
The data lengths used in [26] are not stated. [120] starts input at point 22000 (fig. 15), and uses 1000 points as washout, 3000 points as training, and 1000 points as testing. These values are chosen to ensure both the training and testing phases include both normal sleep and apnea periods.
@@ 8.1.1 Linear Memory Capacity
Certain tasks are interesting because they directly reveal properties of a given Reservoir Computer. One of these is the Linear Memory Capacity task, which quantifies the fading memory of the reservoir. This was introduced by Jaeger [123], and has been investigated in the context of reservoir computing and used as a benchmark task [38, 14, 51, 41, 119].
The Linear Memory Capacity of a given Reservoir Computer that takes scalar input is defined as follows. Consider an input stream $u(t)\in U[-1,1]$. Train the reservoir to reproduce this input, delayed by a number of timesteps $k$: the target output is ${\hat{v}}=u(t-k)$, the observed output is $v_{k}(t)$, the linear memory capacity for this delay is defined as the covariance squared of the delayed input (target output) and the observed output, normalised by the variances of the input and the observed output [123, eqn.13]:
@@
The total Linear Memory Capacity is the sum over all delays [123, eqn.15]:
@@
For higher $k$, each $MC_{k}$ tends to decrease (and the corresponding validation NRMSE tends to increase), which means that inputs after longer delays are remembered less well, and are dominated by noise. In the formal definition of $MC$ (eqn. 22), the sum over delays goes to infinity; however the small values at high $k$ are essentially noise, and should be neglected, so in practice a cutoff is used. Dambre et al. [124, SupMat3.2] define a threshold based on the size of the reported capacity. Dale [45] uses a cutoff of $k_{max}=2N$ (number of nodes).
@@ 8.1.2 Nonlinear Memory Capacity
While the Linear Memory Capacity measure gives a useful quantification of a reservoir’s memory capacity, it cannot on its own model the full computing power of dynamical systems. [124] generalise the definition of linear memory capacity in a way that allows them to define various non-linear capacities, too: they call this information processing capacity (IPC). These measure a Reservoir Computer’s ability to compute a nonlinear function of past inputs, for example, a cubic function of delayed inputs, such as $u^{3}(t-1)$ or $u(t-1)u^{2}(t-2)$. The cases where the polynomials involve different time delays are called cross memory capacities [125].
While [125] considers quadratic polynomials, [124] considers complete sets of orthogonal polynomials, to cover all possible non-linear capacities and cross memory capacities with no double counting. They use normalised Legendre polynomials for inputs $u(t)\in[-1,1]$. They note that different sets of orthogonal polynomials should be used for different input distributions, for example, Hermite polynomials for Gaussian-distributed inputs. Orthogonal sets of trigonometric functions can also be used as the basis.
As for linear memory capacity, the reservoir is trained to output the relevant (here, polynomial) function of its delayed input. For a given degree $d$ of polynomial (linear, quadratic, cubic, etc), the contributions for all delays $k$, including all combinations of delays in the cross memory capacities, are summed. Then the total non-linear memory capacity, or IPC, is the sum over all polynomial degrees. The total IPC, from the contributions of all the linear and non-linear polynomials, is $N$ [124].
@@ 8.2 Rank-based measures
Two particular rank-based measures: kernel quality (kernel rank, KR), and generalisation rank (GR), were first introduced in [126, 127]. These have been adapted for use with ESNs in two different ways in the literature, discussed below. Calculating the rank is discussed in section 8.2.3.
@@ 8.2.1 Kernel Rank
Vidamour et al. [128] use a direct translation of the original definitions, and implement this measure as follows. Consider $S$ maximally distinct input streams, with values drawn from $U[-1,1]$, each of length $T$, and measure the reservoir state at the end of each input stream. KR is the rank of the resulting $N\times S$ matrix $[\mathbf{x}_{1}\mathbf{x}_{2}\ldots\mathbf{x}_{S}]$, see algorithm 1.
Algorithm 1 Vidamour KR [128]
@@
Dale et al. [5] use a different adaptation, which is computationally less intensive (requiring $S$ inputs, rather than $S\times T$), and adapted to time series tasks (where the reservoir state at each timepoint is used), but is further from the original definitions. In this approach, there is a single input stream, of length $S$, and the reservoir state is measured at each timepoint. The stream’s values are again drawn from $U[-1,1]$, thereby making them maximally distinct. See algorithm 2.
Algorithm 2 Dale KR and GR [5]
@@ 8.2.2 Generalisation Rank
Vidamour et al. [128] again use a direct translation of the original definitions. Consider $S$ input streams, each of length $T$, each with values drawn from $U[-1,1]$, except that the last few values in each stream are set to be the same for all streams. Measure the reservoir state at the end of each input stream. GR is the rank of the resulting matrix, see algorithm 3.
Algorithm 3 Vidamour GR [128]
@@
Dale et al. [5] again use a different adaptation of the original definitions. The algorithm for GR is the same as for KR, except that the streams’ values are drawn from a reduced range $U[-0.1,0.1]$, thereby making them similar. See algorithm 2.
@@ 8.2.3 Calculating the rank
An alternative to this integer-valued rank is the real-valued effective rank [129, 130]. Normalise the singular values $\sigma_{i}$ to sum to one: $p_{i}=\sigma_{i}/\sum_{i}\sigma_{i}$. The effective rank is defined as $\exp(-\sum_{i}p_{i}\ln p_{i})$. If $R$ of the singular values are the same, and the rest are zero, then $p_{i}=1/R$ or 0, giving an effective rank of $R$, which is the same as the standard rank value. For other cases, effective rank gives continuous values, has no arbitrary cutoff, and weights the singular values according to their size, potentially giving a more meaningful result.
The number of measured states $S$ should not be less than $N$, since the rank of the resulting matrix is $\leq\min(S,N)$. The measured rank increases with $S$ until it eventually converges [5, sec.D.a]. Preliminary investigation should be performed to establish a suitable value for $S$.
8.2.4 Generalising the rank measures
Given the issues surrounding the definition and comparison of KR and GR in the context of reservoir computing, other measures are being developed to capture similar information, while being better defined. For example, a family of entropy transformation measures, that evaluate how the entropies of the input and output data are related, have been developed [131, 132].
@@ 8.3 Benchmarks for ‘free’
Certain tasks, such as MC and GR, can be calculated from the same data output from the ESN [130]. Both tasks use the same input; the observations of the state can be used to calculate the KR, and an output layer can be trained on the same observations to find the memory capacity. This is useful when using slow physical reservoirs, or computationally intensive search methods such as CHARC.
@@ 8.4 CHARC
Memory capacity, Kernel Rank, and Generalisation Rank are all relevant properties of Reservoir Computers, but the aim should not be to maximise them all. There is no agreed upon principle that allows us to correlate these values of these metric. Indeed, different tasks are best accomplished by reservoirs with different properties [5]. Given this, they construct a framework based on Linear Memory Capacity, Kernel Rank, and Generalisation Rank. The Reservoir can then be evaluated not through its performance at the tasks used to measure these properties, but instead by the breadth of the behaviour space available to the substrate or model it is built upon.
@@
Dale et al. [5] map four of the benchmark tasks described here to their positions in the CHARC behaviour space. One piece of interesting further work might be to map more of them, in order to enable studies to select tasks that not only take a range of different forms, but also correspond to different areas of Reservoir behaviour spaces.
@@ 9 Best practices for benchmarking
Having looked at many of the benchmarks used in the field of Reservoir Computing individually, we now describe some best practices for using them. These can be considered the minimum criteria for reporting results and ensuring reproducibility. However, some of the discussions above demonstrate that meeting these criteria can be a more subtle and involved task than it may appear at first glance.
@@
This section primarily concerns running and reporting benchmarks in reservoir computing. For more general introductions and tutorials on designing, training and using RCs, see [133, 22, 15].
@@ 9.1.1 Using existing datasets
Some equation-based benchmarks, such as the NARMA (section 4.1) and Mackey–Glass (section 5.1) systems, rely on specific parameter values. Different values may lead to different behaviour, as with the Mackey–Glass system, where a $\tau$ of 16.8 or more will lead to chaotic behaviour. Sometimes several different sets of values are used in the literature.
@@ 9.1.2 Introducing new datasets
As noted in [24], if a benchmark is based on a differential equation and numerically integrated, then the values used in the integration, such as timestep and grid size, or the details of a more sophisticated algorithm, can change the detailed generated dataset.
@@
If using a new experimental or generated data set, it needs to be published with the results, and accompanied by metadata. Gebru et al. [66] provide comprehensive guidelines for what metadata should be provided, covering issues including who created the dataset and who funded them, how and when the data was collected or generated, ethical concerns, the structure of the data, sampling, noise and errors, labelling, data cleaning and pre-processing, previous uses, distribution, and maintenance. Gilpin [24]’s catalogue of 131 chaotic dynamics benchmark systems follows the approach given there.
@@ 9.2 Experimental method
The first aspect of experimental setup that needs to be reported are the parameter values of the Reservoir Computer used. One example of this being done well can be found in [1], where the setup for the Mackey–Glass benchmark is described as having 400 nodes, bias input, inserted noise, and output feedback. When reporting an experiment involving an ESN, some parameters that should be reported are:
@@
A guide to setting these parameters can be found in [22].
@@
It is important to record the length of the washout set, particularly for common dataset-based benchmarks, where different washout lengths mean that the Reservoir Computer is being tested and trained on different subsets of the dataset. The training length of the reservoir can have an effect on its performance: up to a certain training length, increasing the training will lead to more consistent results across experiments, leading to a smaller standard deviation of different performances. After this saturation is reached, increasing the training length has little effect, and may result in overfitting. This saturation can be seen both in performance-based tasks such as NARMA, and value-based tasks such as Kernel Rank [45].
@@
Similarly, different testing lengths may affect overall performance. For example, in a free-running prediction task (section 3.1.2), the experiment that uses a longer testing set may yield worse performance, due to the accumulation of errors leading to growing errors [12].